OpenStudy (anonymous):

[INTERGRAL TEST] Find the values of p for which this series is convergent. (See my subsequent post)

OpenStudy (anonymous):

Now I hope somebody will come

OpenStudy (anonymous):

maybe by root test: \[\sqrt[n]{1/Ln( n) ^{np}}=1/Ln (n) ^{p}<1\] so p>1/Ln(n)

OpenStudy (anonymous):

No I have not learned root test yet

OpenStudy (anonymous):

integral test?

OpenStudy (anonymous):

Yes integral test

OpenStudy (anonymous):

Wait, I actually got the answer to this question. There is another problem that asks for the same thing but is slightly more complicated. Let me edit the first post

OpenStudy (anonymous):

ok lets see if we can find an anti derivative maybe try \(u=\ln(x)\) so \(du=\frac{dx}{x}\) and you have \[\int \frac{du}{u^p}=\frac{u^{p+1}}{p+1}=\frac{\ln(x)^{p+1}}{p+1}\]

OpenStudy (anonymous):

Which one are you doing?

OpenStudy (anonymous):

See my latest attachment. That is the one I really need help with

OpenStudy (anonymous):

wow those hits just keep coming are we supposed to find an anti derivative for that thing??

OpenStudy (anonymous):

That is what I was thinking, but then I end up getting stuck when I try to evaluate the integral

OpenStudy (anonymous):